OpenStudy (anonymous):

Find the integral of (e^2x-e^-2x)/(e^2x+e^-2x) ???

OpenStudy (anonymous):

Don't know where to start with this problem... Here's a drawing of it if it's more helpful. 

OpenStudy (fibonaccichick666):

Are you familiar with the exponential forms of the trig functions?

OpenStudy (anonymous):

e^2x+e^-2x=t

OpenStudy (anonymous):

differntiate it first

OpenStudy (fibonaccichick666):

If so, this becomes really easy

OpenStudy (kainui):

Since this is just tanh(2x) you can say the derivative is simply 2sech(2x)!

OpenStudy (anonymous):

Wait what.. I thought the integral of tan was -ln (cosu) +C?

OpenStudy (fibonaccichick666):

wait wait, first let's get to tanh @Kainui

OpenStudy (anonymous):

e^2x+e^-2x=t on differntiating (2*e^2x -2*e^-2x)*dx=dt 2*(e^2x -e^-2x)*dx=dt (e^2x -e^-2x)*dx=dt/2

OpenStudy (fibonaccichick666):

and dixie, not quite the same, if you are not familiar with the exponential form of trig fns. @dixiemitsy , have you seen this before? \(sinx=\frac{e^x-e^-x}{2}\)

OpenStudy (anonymous):

Nope haven't seen it.. and @g17 does that mean you wouldn't have to take the integral of the bottom because it would cancel out?

OpenStudy (fibonaccichick666):

ok then the approach that everyone except g17 is taking, will not make sense to you

OpenStudy (anonymous):

Ok, thanks :P

OpenStudy (fibonaccichick666):

Do you understand g17's approach?

OpenStudy (anonymous):

Yup, got it now :))

OpenStudy (fibonaccichick666):

alrighty good luck! :)

OpenStudy (fibonaccichick666):

Just to make me feel better, what did you get?